### v1

# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Despite extensive investigations, many aspects of charge density waves (CDWs) remain elusive, especially the relative roles of electron-phonon coupling and Fermi surface nesting as the underlying driving mechanisms responsible for the emergence of the CDW vector $\boldsymbol{Q}_{CDW}$. It is puzzling that even though electrons interact strongly with optical phonons in many correlated systems, the actual mode softening is of an acoustic mode. Here we consider monolayer 1H-NbSe2 as an exemplar system, and through an accurate computation of the phonon self-energy, including its off-diagonal components. We provide compelling evidence that the relevant mode is a longitudinal optical phonon that softens by anti-crossing several intervening phonon bands. We also show that $\boldsymbol{Q}_{CDW}$ is fixed by the convolution of the susceptibility and electron-phonon coupling, and that the softened phonons are circularly polarized.
I Introduction
Transition-metal dichalcogenides (TMDCs) are well-known for their layered structures in which transition-metal atoms are sandwiched between two layers of chalcogen, creating a structure that supports a wide variety of electronic interactions driven by valence electron configurations and the spatial symmetries of the unit cell  [TMDC1, TMDC2, TMDC3]. In these compounds, the resulting physical phenomena include the coexistence of charge density wave (CDW) and superconductivity (SC) in NbSe2 [ScCdwExp, ScCdw2]. The simultaneous presence of CDW and SC phases offers a unique platform for exploring effects of complex electron correlations in condensed matter systems.
Among the TMDCs, NbSe2 is distinguished as it exhibits an unusually rich tapestry of properties in both the bulk and monolayer phases. Bulk NbSe2 hosts a CDW with a periodic lattice distortion (PLD) [FengZheng] below $T_{\text{CDW}}\approx 33$ K, along with a coexisting SC phase with $T_{c}\approx 7.2$ K. As the number of layers in a film of NbSe2 is reduced the CDW transition becomes stronger and superconductivity weakens. In the 1H-monolayer limit $T_{\text{CDW}}\approx 145$ K and $T_{c}\approx 1.9$ K [ScCdw3, ScCdw4]. Notably, the layered architecture of TMDCs like NbSe2 also makes them amenable to strain engineering and the construction of heterostructures [strain, hetero1, hetero2]. This adaptability enables a wide range of applications in nanoelectronics, optoelectronics, and quantum devices [Vaskivskyi2015Fast, Akinwande2017Two, Clarke2008Superconducting], where manipulating electronic properties at the nanoscale is of key importance.
The nature of electron-phonon coupling varies across the landscape of materials. In some cases, electron density gets localized and distorts the lattice around the site, effectively enhancing the electronic mass, while in other cases the phonon modulates the bond lengths, strengthening some hoppings and weakening others[polaron1, polaron2]. In polaron physics the former are known as Holstein polarons and the latter as Peierls polarons.[Berciu], with drastic differences in their properties. For Holstein polarons – with increasing electron-phonon coupling, electrons become increasingly more trapped in their polaron clouds and eventually undergo a metal-insulator transition[MIT]. In contrast, for the Peierls polarons, the bonds with strengthened hoppings can link up along certain chains, so that the polaron mass can remain small even for strong electron-phonon coupling. Holstein polarons thus unlikely candidates for bipolaron superconductivity[Mott], but this remains a possibility for Peierls polarons.
This difference should persist in the CDW physics as well. For example, many high-frequency oxygen modes in the cuprates are of Peierls form, modulating copper-to-oxygen hopping, so that the CDW transition would be expected to involve the softening of these optical modes[CuprateSoften1, CuprateSoften2, CuprateSoften3], although this softening has been observed only rarely in experiments.
The softening of an optical phonon to zero energy leads to a form of “quantum strangeness” [Strangeness] in the sense that the soft mode Kohn anomaly[Kohn] must cross all lower-lying phonon modes. Modes of different symmetries can cross without interactions, but those of the same symmetry (irreducible representation (irrep) of the dynamical matrix), interact to open an energy gap or ‘anticross’, where eigenfunctions interchange their characters, see Figure 1(a). At critical softening the Kohn anomaly thus resembles a “Kohn ladder” or a Kohn anomaly which is broken up into many segments separated by anticrossings at ascending energies.
Such a ladder has been predicted[Gutzcharge] and observed[Ding] in the cuprates, where the resulting CDW is consistent with a known optical phonon symmetry of the parent compound[Forgan]. In NbSe2, an experimental study examined but ultimately rejected the possibility of a Kohn ladder[Rez]. However, in monolayer 1H-NbSe2, which lacks inversion symmetry, two possible CDW distortions have been reported, which are consistent with breathing-in and breathing-out distortions of an optical phonon[NbSe2]. Here we re-explore this issue by examining the contribution of electron-phonon coupling to the phonon self-energy to uncover what controls the CDW $q$-vector.
II NbSe2
NbSe2 is a canonical metallic superconducting CDW material, with the superconducting state occurring at a temperature $T_{c}\ll T_{\text{CDW}}$ [Wilson1975, Moncton1977]. Near $T_{\text{CDW}}$, there is a strong Kohn anomaly associated with a soft phonon and a static CDW exhibiting a $3\times 3$ PLD in real space [EPH3, EPH4]. This corresponds to a triple CDW wavevector $\mathbf{Q}_{\text{CDW}}\approx\frac{2}{3}\overline{\Gamma M}$ [Soumyanarayanan2013, EPH5].
Unlike NbS2, NbSe2 does not exhibit magnetic response and it is not close to any magnetic instability in its CDW phase [magnetic, magnet2]. The electronic spin-orbit energy splitting at the Fermi level is negligibly small compared to the CDW energy scale [Johannes2006]. The Fermi surface of NbSe2 is strongly anisotropic, composed of pockets around the $\Gamma$ and $K$ points, preventing perfect nesting at the Fermi energy [Flick, FengZheng, FermiGap]. These features of the electronic structure imply that the PLD-CDW instability is far from conventional being and that it results from a combination of strong electron-phonon coupling and Fermi surface topology [topo1].
Survival of the superconducting state in the presence of such a strong CDW transition indicates that the CDW and superconducting regions avoid each other, confirming a strongly anisotropic Fermi surface within the CDW state [new_nbse2, 2new_nbse2]. This unconventional picture is distinct from the commonly observed single-acoustic-mode softening accompanied by the appearance of a uniform CDW gap around the Fermi surface. Indeed, density functional theory (DFT) based calculations show that both longitudinal acoustic and longitudinal optical phonon modes have strong coupling with electrons on the Fermi surface, and that spin-orbit effects do not play a major role.
However, many questions about the origin of the CDW in NbSe2 remain controversial  [Johannes2006, Borisenko2009, Kiss2007, Calandra2009, Plummer]. In this study, we will focus on the $T_{c}\ll T_{\text{CDW}}\leq T$ regime, and employ DFT-based calculations to address two questions: How does the optical-acoustic phonon mode mixing relate to the Kohn anomaly? And is $\mathbf{Q}_{\text{CDW}}$ determined by the peak in the charge susceptibility or by the anisotropic electron-phonon coupling?
Figure 1: (color online) Phonon dispersions in NbSe2. (a) $\sigma$-dependent phonon spectrum. (b) $\sigma$-dependent self-energy spectrum for acoustical bands. (c) $\sigma$-dependent self-energy spectrum for optical bands. (d) Variation of unit cell parameters vs $\sigma$. Blue line gives the change in the lattice parameter, while red line gives the distance between Nb and Se atoms with respect to the change in $\sigma$.
III Calculating electron-phonon coupling
In conventional many-body approaches, the bare phonon Green’s function is calculated without electron-phonon interactions, which are incorporated subsequently in the phonon self-energy [Mahan]. In contrast, DFT calculations of the phonon spectra do not make an explicit separation into bare and dressed phonons, which complicates comparisons between various phonon results  [Giustino2017, Marini2001, Heid2010, Harkonen2020].
We first use Quantum Espresso (QE) [QE_refs1, QE_refs2] for DFT-based structural optimization and electronic structure computations, see Appendix D for details. A Fermi-Dirac type smearing $\sigma$ is invoked to simulate thermal effects and define an effective temperature $T_{\sigma}=\sigma/k_{B}$, where $k_{B}$ is Boltzmann’s constant. We then use density-functional perturbation theory (DFPT), as implemented in the QE software[QuantEspRenorm], to directly phonon frequencies as a function of $\sigma$. Electron-phonon coupling matrix elements are then computed by using the EPW package [EPW]. The original EPW code gives the absolute value of the electron-phonon coupling and the diagonal elements of the self-energy matrix. To assess the couplings between the phonon modes and the phonon self-energy, we modified the EPW code[EPWmodified] to calculate the off-diagonal elements $\lambda\neq\lambda^{\prime}$ of the phonon self-energy[typy] and the coupled phonon eigenstates.
Figure 2: (color online) Decomposition of self-energy to determine the origin of the $q$-vector close to the instability temperature, $\sigma=0.00565$ Ry. Top row (left to right): (a) Phonon dispersions for various modes, numbered with different colors. (b) Susceptibility, $\chi(q)$ (dashed red line, right-axis)), and $k$-averaged electron-phonon coupling $G(q)$ (remaining curves, left axis). (c) $G(q)\chi(q)$. (d) Diagonal self-energy $\Pi^{\prime}_{\lambda\lambda}$. (e,f) Real (e) and imaginary (f) parts of the off-diagonal self-energy $\Pi^{\prime}_{0\lambda}$, where 0 denotes the soft-mode phonon mode (blue). (g) $1/\Omega$. and (h) $\Pi$. Position of the soft-mode $\bf Q_{CDW}$-vector (solid black vertical lines) and $\bf q=2/3\,\overrightarrow{\Gamma M}$ (dashed black vertical lines) are shown.
Figure 2 presents the phonon spectrum along the $\Gamma-M$ in panel (a), while the remaining frames provide results for various components of the calculations to gain insight into the nature and origin of the CDW instability. Modes 3, 6, and 8 have predominantly longitudinal symmetry and interact strongly with each other. Modes 2 and 7 are predominantly transverse, but since the modes are strictly longitudinal or transverse only along the high-symmetry lines such as $\Gamma-M$, these longitudinal and transverse modes interact away from the high-symmetry lines. These are the only modes with strong electron-phonon coupling, and therefore, we will we focus mainly on these modes. The sharp dip in the phonon dispersion of mode 3 in panel (a) is the Kohn anomaly that drives the CDW instability[Kohn].
Figure 1 summarizes the evolution of the phonon spectra (a-c) and lattice parameters (d) for various $\sigma$ values. Strong $T$-dependence of the phonon frequencies and the appearance of phonon softening are clear evidence that the interactions are inseparably incorporated in the DFT calculations. Strong $T$-dependence of the atomic positions and lattice constants in the numerical results shown in Figure 1(d) suggest the importance of multiphonon processes (noninteracting phonons do not involve thermal lattice expansion). The the phonon self-energy is calculated non-perturbatively here in terms of dressed phonons, which leaves open the questions: How does one extract the bare phonon frequencies? And is the DFT based self-energy consistent with the field theory renormalization of the bare phonon frequencies?
To reconstruct the dynamical matrix in the presence of electron-phonon coupling, we start with a general framework where multiple phonon modes interact with multiple electron bands in the vicinity of the Fermi energy. [Mode and band are used in the Born-Oppenheimer sense where the ionic and electronic degrees of freedom are decoupled.] The electron-phonon interaction is handled perturbatively, see Appendix B for details.
A phonon self-energy matrix $\Pi_{\lambda\lambda^{\prime}}$ is introduced which includes off-diagonal coupling between the phonon modes $\lambda$ and $\lambda^{\prime}$ as well as the electron-phonon vertex function $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$; here the electron state of wavevector $\boldsymbol{k}$ in band $n$ interacts with the electron state $\boldsymbol{k}+\boldsymbol{q}$ in band $m$ by exchanging a phonon with wavevector $\boldsymbol{q}$ in the mode $\lambda$. Since NbSe2 only has one band crossing the Fermi energy, we take $n=m=1$, and remove the electron band index for simplicity. Spin-orbit coupling, which is quite small in NbSe2, is ignored.
In the multi-mode phonon basis[Giustino1], the renormalized phonons are given by the poles of the interacting Green’s function:
$\det\Bigl{\{}(\Omega^{2}-{\boldsymbol{\omega}_{\boldsymbol{q}}^{2}})\operatorname{\boldsymbol{I}}-2\,{\boldsymbol{\omega}_{\boldsymbol{q}}}\,{\boldsymbol{\Pi}}(\boldsymbol{q},\omega_{q})\Bigr{\}}=0$ (1)
where $\boldsymbol{\omega}_{\boldsymbol{q}}$ is a diagonal matrix with elements $\omega_{\lambda,\boldsymbol{q}}$ in the phonon eigen-mode space describing the bare-phonon frequencies and ${\boldsymbol{\Pi}}(\boldsymbol{q},\omega)$ is the self-consistent electron-phonon self-energy given by,
$\Pi_{\lambda,\lambda^{\prime}}(\boldsymbol{q},\omega)$ $=$ $\sum_{n,m,\boldsymbol{k},\sigma}\,G_{\lambda,\lambda^{\prime}}(n,\boldsymbol{k};m,\boldsymbol{k}+\boldsymbol{q})$ (2)
$\times$ $~{}~{}~{}~{}~{}~{}\frac{f(\xi_{m,\boldsymbol{k}+\boldsymbol{q},\sigma})-f(\xi_{n,\boldsymbol{k},\sigma})}{\xi_{m,\boldsymbol{k}+\boldsymbol{q},\sigma}-\xi_{n,\boldsymbol{k},\sigma}+\omega}$
with $\xi_{n,\boldsymbol{k},\sigma}$ describing the electron energy in band $n$, measured from the Fermi level and $f(x)$ is the Fermi-Dirac distribution. Here
$G_{\lambda,\lambda^{\prime}}(n,\boldsymbol{k};m,\boldsymbol{k}+\boldsymbol{q})=\mathfrak{g}_{n,m,\lambda}(\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q})\,\mathfrak{g}_{m,n,\lambda^{\prime}}(\boldsymbol{k}+\boldsymbol{q},\boldsymbol{k})$
where
$\mathfrak{g}_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ $=$ $\Bigl{(}\frac{\hbar}{2M\omega_{\lambda,\boldsymbol{q}}}\Bigr{)}^{1/2}g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})~{}~{}~{}~{}~{}~{}$ (4)
and $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ is the microscopic complex electron-phonon coupling given by[Giustino1]
$g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})=\langle|\Psi_{\boldsymbol{k}+\boldsymbol{q},m}|\nabla_{\boldsymbol{q}}V_{\boldsymbol{q}}.\boldsymbol{\xi}_{\lambda\boldsymbol{q}}|\Psi_{\boldsymbol{k},n}\rangle.$ (5)
where $|\Psi_{\boldsymbol{k},n}\rangle$ are the electron Bloch states, see Appendix B for details. Inhomogeneous $\boldsymbol{k}$ dependence in Eq.(5) and hence in Eq.(3) is the result of the anisotropy and the sensitivity of the Bloch wave functions to the crystal structure. In NbSe2 the relevant sections of the Fermi surface responsible for the strong electron interaction are in the K-pockets near the zone boundaries.
Note that $\Pi_{\lambda\lambda^{\prime}}$ depends on the bare phonon frequencies, $\omega_{\lambda\boldsymbol{q}}$, but that an exact solution for the dressed phonons can be obtained in principle by replacing self-consistently the bare terms by the dressed terms [Hedin, Mahan]. Below, we will refer to the replacement of $\omega_{\lambda\boldsymbol{q}}$ by $\Omega_{\lambda\boldsymbol{q}}$ as the Hedin model.
We now discuss how the $\boldsymbol{k}$ dependence of $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ drives an unconventional phonon softening and CDW formation distinct from the Peierls mechanism[gruner1994density].
In the limit of negligibly weak lattice potential, the Bloch states can be approximated by plane waves, so that $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ in Eq.(5) becomes dependent only on the momentum exchanged with the phonons[Grimvall]. The phonon self-energy in Eq.(2) can then be written as:
$\Pi_{\lambda,\lambda^{\prime}}(\boldsymbol{q},\omega)$ $=$ $G_{\lambda,\lambda^{\prime}}(\boldsymbol{q})\chi_{0}(\boldsymbol{q},\omega)$ (6)
The second term on the right hand side of Eq. 6 is closely related to the Lindhard electronic susceptibility (dropping the band index n and considering the spin degeneracy factor 2),
$\chi_{0}(\boldsymbol{q},\omega)=2\,\sum_{\boldsymbol{k}}\,\frac{f(\xi_{\boldsymbol{k}+\boldsymbol{q}})-f(\xi_{\boldsymbol{k}})}{\xi_{\boldsymbol{k}+\boldsymbol{q}}-\xi_{\boldsymbol{k}}+\omega+i\delta},$ (7)
which we calculate using Maximally Localized Wannier Functions (MLWFs) [Wannier]. We further simplify the computation by taking the adiabatic limit ($\omega\to 0$). [Non-adiabatic corrections in NbSe2 are found to be small.] Effects of the $\boldsymbol{k}$ dependence of $g_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ can be seen in $\Pi_{\lambda,\lambda^{\prime}}$ evaluated using the limiting cases given by Eqs. (2) and (6).
Figure 2(a) shows the DFPT phonon dispersions[Baroni] calculated at a temperature slightly above $T_{CDW}$. The CDW transition is accurately predicted at the wavevector $\boldsymbol{Q}_{CDW}\simeq 2\,{\overrightarrow{\Gamma M}}/3$ which agrees with the experimental value, see Appendix A for details. It is convenient to introduce a new quantity here:
$\Pi^{\prime}_{\lambda\lambda^{\prime}}(\boldsymbol{q},\omega)=\sqrt{\omega_{\lambda\boldsymbol{q}}\omega_{\lambda^{\prime}\boldsymbol{q}}}~{}\Pi_{\lambda\lambda^{\prime}}(\boldsymbol{q},\omega),$ (8)
so that Eq.(2) can be recast as
$\begin{split}&\Pi^{\prime}_{\lambda,\lambda^{\prime}}(\boldsymbol{q},\omega)=2\frac{\hbar}{2M}\sum_{n,m,\boldsymbol{k}}\frac{f(\xi_{m,\boldsymbol{k}+\boldsymbol{q}})-f(\xi_{n,\boldsymbol{k}})}{\xi_{m,\boldsymbol{k}+\boldsymbol{q}}-\xi_{n,\boldsymbol{k}}+\omega}\\ &\times~{}~{}~{}~{}~{}~{}\,g_{n,m,\lambda}(\boldsymbol{k},\boldsymbol{k}+\boldsymbol{q})\,g_{m,n,\lambda^{\prime}}(\boldsymbol{k}+\boldsymbol{q},\boldsymbol{k}).\end{split}$ (9)
We will refer to $\Pi^{\prime}$ as an auxilary self-energy since it has dimension (energy)2. Note that the overall phase associated with gauge non-invariance of $\mathfrak{g}_{n,m,\lambda}(\boldsymbol{k};\boldsymbol{k}+\boldsymbol{q})$ does not affect its off-diagonal elements.
To understand the origin of phonon softening, we examine ab initio and many-body results based on Eq.’s(2) and (6). This requires a modification of the QE+EPW results, since the DFPT already incorporates the main phonon renormalizations (Appendix B). In order to calculate $\Pi^{\prime}$ in Eq.(8), we must invert the dynamical matrix to get a handle on the nature of the bare phonons. This involves two issues. (i) Ambiguity in inverting the dynamical matrix, which we resolve in Section VII and Appendix C. And (ii) We need the off-diagonal self-energy elements, which are essential for the adiabatic term, even though these are unimportant for the non-adiabatic corrections, extraction of off-diagonal self-energies elements is discussed in Appendix A.
IV What determines $\boldsymbol{Q}_{CDW}$?
The debate here has been on whether $\boldsymbol{Q}_{CDW}$ is dominated by the susceptibility in Eq.(7) or the electron-phonon matrix elements in Eq.(3). However there is another potential mechanism. In Eq. 4 , the pre-factor contains $1/{\omega_{\boldsymbol{q}}}$ in $|\mathfrak{g}|^{2}$. This, following Hedin[Hedin], can be replaced by $1/{\Omega_{\boldsymbol{q}}}$, which acts as a Stoner denominator and causes the diagonal element $\Pi_{\lambda\lambda}$ to diverge when $\Omega_{\bf q}\to 0$ for $\boldsymbol{q}=\boldsymbol{Q}_{CDW}$ as $T\to T_{CDW}$. A similar argument has been invoked for the CDW physics in the cuprates.[Gutzcharge]
Fig. 2 presents a decomposition of the self-energy of all phonon modes in panel (a) along the $\Gamma\rightarrow M$ symmetry line. We first focus on the soft-mode mode (dark green lines). $\Pi$ (panel h) shows a sharp, nearly-divergent peak at the soft-mode $\boldsymbol{q}$-vector, which arises entirely from the $1/\Omega_{q}$-factor (panel g). In contrast, $\Pi^{\prime}=\omega\Pi$ (panel d) has a weak peak at the same $\boldsymbol{q}$-vector. This sharpening of a weak peak is a typical effect of a Stoner denominator. If the electron-phonon matrix element were $\boldsymbol{q}$-independent, then $\Pi^{\prime}$ would peak at the same $\boldsymbol{q}$-value as the susceptibility $\chi$ (dashed red curve in panel b), which is not the case. Similarly, if $\chi$ were constant, the peak would follow $G$ (green curve in panel b), but that is also not the case. The product $G\chi$ (panel c) comes closer, indicating that both $\chi$ and $G$ possess significant $\boldsymbol{q}$ and $\boldsymbol{k}$ dependence, such that the full integral over $\boldsymbol{k}$ correctly reproduces the self-energy peak in panel d. The peaks in both $\Pi$ and $\Pi^{\prime}$ (solid-black vertical lines) are in reasonable agreement with the experimental value $\boldsymbol{Q}_{CDW}\simeq 2~{}{\overrightarrow{\Gamma M}}/3$ (black vertical dashed line). Note however that the off-diagonal terms in $\Pi^{\prime}$ [panel f] due to the phonon mode-mode coupling can lead to a small additional shift of the $\boldsymbol{Q}_{CDW}$.
When we extend preceding discussion to encompass the the self-energy in the full Brillouin zone, we see that (Fig. 1 (b)) the largest self-energy peak falls along the $M\rightarrow K$-line. This sharp peak is associated with a weak nesting on the Fermi surface, see Appendix A for details. Insight can be gained by recalling that in many superconductors, the superconducting gap can be captured in the optical spectra only in the dirty-limit where momentum conservation breaks down[MatBar]. A similar mechanism may be at play with the second peak along the $M\rightarrow K$-line in our case and its absence in experiments. Note that Figure 1(b) is for the clean limit. To capture the dirty limit, however one must incorporate the breaking of the momentum conservation due to random impurities, which requires averaging $G_{\lambda,\lambda^{\prime}}$ over $\boldsymbol{k}$, leading to the form of Eq. 6. Fig. A.2 (b) (Appendix A) shows that the averaging procedure smears out this weak nesting feature while preserving the peak along $\Gamma M$ with a modest loss of intensity. A similar conclusion was reached in Ref. Plummer.
V Coupling of bare phonon bands
To extract the bare phonons, the divergent $1/\Omega_{\lambda,\boldsymbol{q}}$-factor in Fig. 2(g) must be replaced by $1/\omega_{\lambda,\boldsymbol{q}}$ (Appendix C). Once this is done, we can go beyond the above single mode picture by analyzing the $T$-dependence of the diagonal self-energies in Figs. 1 (b) and (c). We discuss the resulting anticrossing phenomena by considering two distinct cases below.
V.1 Longitudinal-transverse anticrossing
Figure 1 (a) shows that reducing the temperature (right to left) results in crossing of the modes 2 and 3 and also of modes 6 and 7 along the $\Gamma-M$ line (see Fig. 2 for notation of phonon modes). Note that the optical modes 6 and 7 display at low temperatures an unusual Mobius topology along the $\Gamma\to M\to K\to\Gamma$ line: if one follows the dispersion of mode 6, starting from the left-hand side $\Gamma$ point, one ends up on mode 7 on the right-hand side. Repeating this process into the next $\Gamma$-point by using symmetry to jump to the left-hand side $\Gamma$-point, and starting on mode 7, we see that only after winding twice about this path do we return to the original point. This behavior can be understood by noting that the longitudinal and transverse modes are orthogonal, so that we normally expect them to cross. However, in an anisotropic material, eigenmodes are strictly longitudinal/transverse only along the high symmetry lines but this is not be the case along a general direction in Brillouin zone, so that the bands then must anti-cross and exchange wavefunction character. Since modes 6 and 7 can only cross along $\Gamma-M$, they must wind twice to go from $6\rightarrow 7\rightarrow 6$ before returning to the starting mode.
A similar phenomenon is at play for modes 2 and 3 in Fig.1, although it is hard to see the crossing between these modes because they are acoustic modes with vanishing frequencies at $\Gamma$. However, the most dispersive mode near $\Gamma$ is longitudinal (L, numbered 3) at the left-hand side $\Gamma$-point. The resulting mode is numbered 2 after the $\Gamma\to M\to K\to\Gamma$ loop is closed. Similar winding numbers have been observed in the cuprates under pressure[Joh], see also Ref. Zak. At higher temperatures (right column in Fig.1) both anti-crossings disappear, suggesting that they are related to phonon softening and electron-phonon interactions. Notably, crossing of bands 6 and 7 is symmetry protected only along the $\Gamma M$ line, where the crossing point forms the tip of a phononic Dirac dispersion. Since this band crossing is driven by phonon softening, it constitutes an electron-phonon-coupling induced topological phase transition.
Insight into the mode crossing of modes 6 and 7 is also obtained by looking at the self-energies. As $T$ increases, the longitudinal mode 7 hardens and stops overlapping the transverse mode. At the crossing point, a part of the self-energy, especially along $K$ to $\Gamma$ jumps from the lower to the higher (now longitudinal) mode. [The strong spectral weight is always associated with the upper modes near $K$, but due to the change of connectivity, the labels of the modes change.] Effects of anticrossing can, however, still be seen in Figs. 1(b,c), where the spectral weight of the upper mode decreases substantially below the anticrossing temperature, i.e., as $T$ is lowered, the modes start to anticross, and more longitudinal character transfers to the lower modes, along with the strong electron-phonon coupling.
Turning to modes 2 and 3, the crossing along $\Gamma M$ is seen to disappear at essentially the same temperature as for modes 6 and 7, and there is also a transfer of self-energy along $K-\Gamma$, but the effect is less pronounced because the modes do not actually cross.
V.2 Longitudinal-longitudinal anticrossing
Here, the anticrossing of the three longitudinal modes is harder to see in the dispersions as there are no symmetry-allowed crossing. Kohn-anomaly related anticrossings may thus be taken as a demanding test of a phonon code. Although we have shown that QE captures mode-crossing effects for modes 6 and 7, let us now consider modes 3 and 8, which are the lowest and highest longitudinal modes. As expected, the self-energy of mode 3 decreases with increasing $T$. However, the self-energy of mode 8 increases with increasing $T$, and as Figs. 1(b,c) indicate, this is due to transfer of spectral weight from mode 3 (note the anomalous peak along the $MK$ line marked by arrow in panel (b)), which shifts to mode 8 (arrow in panel (c)) with increasing temperature. These results unambiguously demonstrate that the mode softening starts in the highest longitudinal mode, and as temperature decreases the spectral weight transfers to successively lower longitudinal modes via mode coupling, with further spectral weight transfer between nominally longitudinal and transverse modes, since they are not strictly orthogonal at arbitrary $\boldsymbol{k}$-points. In the remainder of this paper, we explore the dynamical matrix to try to extract a coherent picture in terms of ‘bare’ phonon frequencies.
Figure 3 provides further insight by plotting the diagonal self-energy weight on top of the phonon dispersion data from Fig. 1, with results for the clean limit given in the top row (a) and the dirty limit in the bottom row (b). Transfer of spectral weight from high to low frequencies as temperature decreases can be seen, which is somewhat clearer in the dirty limit (bottom row).
Figure 3: (color online) Phonon self-energies. From left to right 0.0057, 0.009, 0.02 eV smearing. From top to bottom: (a) clean-limit and (b) dirty limit; projected in color on phonon bands (color codes from dark to light map into weak to strong phonon self-energy).
VI Dynamical matrix and bare phonon frequencies
We can obtain the bare frequencies $\omega_{\bf q}(T)$ at various temperatures by inverting the dynamical matrix. We would expect the bare frequencies to be approximately independent of $T$, but that is not the case, see Appendix B for details. Note that phonon anharmonicity[anharm] is implicitly included in the first-principles DFPT calculations.)
We must first settle on the form of the dynamical matrix. In Appendix C, we show that, despite some ambiguity, the appropriate choice is:
$\det\Bigl{\{}(\omega^{2}-{\boldsymbol{\Omega}_{\boldsymbol{q}}^{2}})\operatorname{\boldsymbol{I}}+2\,{\boldsymbol{\Omega}_{\boldsymbol{q}}}\,{\boldsymbol{\Pi}}(\boldsymbol{q},\Omega_{q})\Bigr{\}}=0,$ (10)
which provides a conventional matrix equation for the bare phonon frequencies in terms of the dressed ones. Fig. 4 shows the bare frequencies (solid lines) obtained by neglecting the off-diagonal terms in Eq. 10. Although phonon softening can be seen, the mode anticrossings, which are controlled by the off-diagonal terms are absent.
Figure 4: (color online) Single mode calculations for the three longitudinal modes (modes 3, 6, and 8). Solution of Eq. 20 for $\sigma$ =0.0565 Ry, comparing the dressed ($\Omega_{q}$, dashed lines) and bare frequencies ($\omega_{q}$, solid lines).
To understand why $\Omega_{\bf q}$ becomes soft, we rewrite the single mode solution of Eq. 1 as $\Omega_{\bf q}^{2}=\omega_{\bf q}^{2}(1-S_{\bf q})$, where $S_{\bf q}=-2\Pi^{\prime}(\bf q,\Omega_{\bf q})/\omega_{\bf q}^{2}$ is an effective Stoner factor[diSalvo1977]. The instability (soft phonon) occurs at that $\bf q$ for which $S_{\bf q}\rightarrow 1$ first. At $\bf Q_{CDW}$, the bare frequency is $\omega_{\bf Q_{CDW}}=\sqrt{-2\Pi^{\prime}(\bf Q_{CDW})}$. If $\omega_{\bf q}$ depends weakly on $\bf q$, then the soft mode occurs when $|\Pi^{\prime}(\bf q)|$ is at its maximum.
VI.1 Multimode calculations
Figure 5: (color online) (a) Dressed (blue dashed lines) and bare (red solid lines) phonons for $\sigma$ = 0.02 Ry smearing. (b) Same as (a) but for $\sigma$ = 0.0057 Ry. (c) Temperature dependence of the bare phonons for $\sigma$ = 0.02 Ry (orange solid lines) and 0.0057 Ry (green dashed lines). (d) Temperature dependence of the dressed phonons, with colors representing a various phonon modes: 2 (orange), 3 (red), 6 (magenta), 7 (green), and 8 (blue). $\sigma$ = 0.0057 Ry (dashed lines) and 0.02 eV (dotted lines), while the solid lines ($\sigma$ = 0.02 Ry bare phonons) are our best estimate of the high-$T$ limit.
Figs. 5(a) and (b) compare the dressed and bare phonon modes at two temperatures ($\sigma$ = 0.02 and 0.0057 Ry) obtained by solving the full $9\times 9$ dynamical matrix. Modes 1, 4, 5, and 9 have no electron-phonon coupling, so that $\omega_{\bf q}=\Omega_{\bf q}$ at eac temperature, although these modes still exhibit a weak $T$-dependence, panel (c). Other modes display stronger $T$-dependence, especially the longitudinal modes. We will see below that most of the $T$ dependence of the transverse modes is associated with anticrossing related weight transfer from longitudinal to transverse modes.
Figure 5(d) provides evidence of a Kohn ladder related to the softening of modes 8, 7, and 3. The transverse modes have a weaker $T$-dependence, mainly in the $q$-regions where the self-energy shows transfer of spectral weight from longitudinal bands, and both bands show regions of phonon hardening at low temperatures. Notably, mode 7 shows strong softening at higher $T$s, which is not as strong when it interacts with mode 6, where mode 3 starts to soften more rapidly.
VII A Model of a Kohn ladder
We consider a model calculation in which the temperature dependent self-energy is a scaled version of $\Pi_{33}$ in Fig. 2(b). To simulate anticrossing, we assume that the self-energy is initially concentrated entirely on the highest mode, $\Pi_{0,88}=\alpha\Pi_{33}$, where $\Pi_{0}$ is the bare self-energy and $\alpha$ increases as $T$ decreases. Figure 6 presents illustrative results, which are based on taking the bare phonons from the one-band model solutions for $\sigma=0.00565$ Ry. The results illustrate the anticrossing phenomena, and are in general accord with the QE results (green dashed line in (b)). Figure 6(a) uses the clean-limit $\Pi_{33}$, showing that the anomalous nesting peak spoils the agreement with experiment, as expected, while the dirty limit result, Fig. 6(b), restores the good agreement. Fig. 6(b) also show that for $\alpha\leq 5$, most of the softening is in the LO modes 6 and 8, while the LA mode barely changes. For larger $\alpha$, most of the softening is in the LA mode. The latter regime resembles the QE data for $\sigma\leq 0.03$ Ry.
Figure 6: (color online) Model calculations of the evolution of the dressed frequencies at a series of temperatures scaled by $\alpha$. For $\alpha=0$ (highest $T$): one-band solutions for $\omega_{q}$ calculated from $\sigma=0.03$ Ry (black dotted lines); adding off-diagonal $\Pi$, with $|\Pi_{ij}|^{2}=20$ (blue lines). Lower temperatures: $\alpha$ =5 (violet lines), 8 (red lines), 9 (brown lines). Other curves: green dashed line = band 3 from $\sigma=0.07$ Ry data; thick dark red line = band 8 for $\alpha=9$ but $\Pi_{ij}=0$.
Notably, if we turn off the mode-coupling, the softening of mode 8 at $\alpha=9$ (thick dark red curve) is too weak for a CDW instability. This can be understood as follows. Considering two modes only, the anti-crossing results in the renormalized frequencies given by:
$\Omega^{2}_{\pm}=\frac{1}{2}\,\Bigl{[}\Omega_{1}^{2}+\Omega_{2}^{2}\pm\sqrt{(\Omega_{1}^{2}-\Omega_{2}^{2})^{2}+16\omega_{1}\omega_{2}|\Pi_{12}|^{2}}\Bigr{]}$ (11)
where the $\Omega_{1,2}$ are the renormalized frequencies found from Eqs.  1 (without mode mixing) and $\Omega_{\pm}$ include mode mixing. Clearly, $\Omega^{2}_{-}$ can go soft completely at the critical mode mixing threshold $\Pi_{12}^{(c)}$ where
$|\Pi_{12}^{(c)}|^{2}=\frac{\Omega_{1}^{2}\,\Omega_{2}^{2}}{4\omega_{1}\,\omega_{2}}$ (12)
which takes place before $\Omega_{1}^{2}$ goes soft (without mode-mode coupling). Off-diagonal self-energies thus generally enhance $T_{CDW}$. Mode-mode mixing can also induce a shift in $\boldsymbol{Q}_{CDW}$. These effects can be shown for NbSe2 by investigating the difference between $\Omega_{\pm}$ and $\Omega_{1,2}$. Note that while the diagonal self-energies of Eq. 1 are real, the off-diagonal terms can be complex. In the irreducible representation with dimension 3, the Hamiltonian determinant contains the term $-16\omega_{1}\omega_{2}\omega_{3}Re(\Pi_{12}\Pi_{23}\Pi_{31})$. That is, the phonon energies depend on the relative phases between the off-diagonal components of the self-energy matrix. In contrast, for irreducible representations of lower dimension, the off-diagonal term only enters as $|\Pi_{12}|^{2}$.
Interestingly, the initial electron-phonon interaction at high temperatures is longitudinal, so that the main phonon softening predominantly affects longitudinal modes. However, due to anticrossing, the phonon that softens will appear to be from a transverse mode, unless, as in NbSe2, the softening is along a high-symmetry line or point where longitudinal and transverse modes are strictly orthogonal.
VIII Discussion
VIII.1 Giant Kohn anomalies
The approach discussed in this paper would be applicable for understanding giant Kohn anomalies segmented into ladders via mode anticrossings in many correlated $d-$ and flimit-$f-$electron materials. Not including effects of anticrossing cloud underestimate the strength of electron-phonon coupling. Even in graphene, DFT calculations fail to properly capture the Kohn anomaly[graphene].
A study of spin-lattice relaxation argues that most phonons play no role in relaxation, which is dominated by phonons with rotational or intramolecular vibrational patterns.[SLR1] Moreover, even these latter phonons lie at high energies, and onlt the small tail of these phonons at low energies dominates the relaxation. Our study suggests that phonons involved in relaxation are all optical, and that the low energy tail is produced by strong electron-phonon interaction in the form of Kohn anomalies, which push some modes to lower energies (soft modes). A chiral Kohn anomaly has indeed been reported recently in a Weyl semimetal.[Mingda]
Figure 7: (color online) K-point phonon displacements. (a) mode 6, showing three phases of displacement in one cycle; (b) mode 7, and (c) mode 2 showing a single phase; and (d) phonon dispersions with red dots showing $k$-points used, from top to bottom mode 6 (a), 7 (b), and 2(c). Yellow boxes highlight one triangle, while violet boxes highlight one hexagon.
VIII.2 Visualizing soft phonons
By using the phonon visualizer software [phonon_visualizer], we find that the phonon modes in NbSe2 with significant electron-phonon couplings undergo displacements that are circularly or elliptically polarized, raising interesting question as to what happens when a circular phonon softens. For a linearly polarized phonon, the frozen phonon CDW will have the maximum distortion associated with the soft phonon. But circular phonons have both amplitude and phase as variables, and the phase variable could lock in uniformly, adopt spatially random fixed directions (as in order-disorder transitions[OD]), or remain free as in a time crystal[time1, time2]. There is also a striking similarity between the circular phonon displacements and the dynamic Jahn-Teller (dJT) phenomena, which have been suggested as a source for time crystals.
We consider two $q$-values: $\bf q=\overrightarrow{\Gamma K}$, which lies in the region where the longitudinal-transverse anticrossing occurs and $\bf q=2/3\overrightarrow{\Gamma M}$, where the soft phonon lies. These two $q$ points are special in that they are 1/3 of a combination of reciprocal lattice vectors. Remarkably, most of the $K$-point phonons are associated with recognizable dJT displacements. One of the earliest dJT systems involves triangular molecules, such as Na3, which can have a Berry phase that leads to an extra orbital angular momentum with half-integer quantization (a ‘time molecule’).[LongHi, Na3] Similar phenomena can be found on impurity sites in other crystals.
A layer of NbSe2 can be viewed as a lattice of centered triangular molecules, either triangles of Se2 molecules with a central Nb or of Nb molecules with a central Se2. For several $K$-point circular phonons one cycle of rotation amounts to the central atom displacing sequentially toward each of the triangular atoms, leading to a time-averaged symmetrical pattern (yellow squares in Fig. 7), as in Refs. LongHi, Na3. However, a lattice of dJT molecules has extra degrees of freedom, which also have a dJT interpretation. On each of the three steps in a cycle, a pattern forms on one set of atoms, then dissolves and reforms on a second and then a third set before it repeats in the next cycle. We have found three such patterns, enclosed by purple squares in Fig. 7: (i) ‘breathing in’ (panel (b)), where three Se2s move close to a central Nb; (ii) ‘breathing out’ (panel (c)), where the three Se2s move close to an empty site; and (iii) Kekule-distortion of a hexagon, where panel (a) illustrates the three phases of a cycle. It has been suggested that Pauling’s resonating valence bond (RVB) of benzene could be a form of dJT effect,[Pauling]. Here we have a circular phonon form of this effect.
Figure 7 illustrates displacement patterns for several $K$-point phonons. Panel (a) shows the Kekule-like pattern in a hexagon of alternating Nb and Se2 pairs (violet box) in the three steps involved in one cycle of mode 6.111This is not strictly a Kekule pattern, as a cycle contains three phases, with alternating short bonds 12-34-56, 23-45-61, and an undistorted hexagon. To confirm our identification of this pattern, we show in Appendix E that a similar pattern arises in graphene, which is a lattice of benzene rings.. Panel (b) shows the breathing-in mode in mode 7, while panel (c) depicts the breathing-out mode of mode 2. The Kekule state is involved in the softening between modes 6 and 7, but it is not the ground state.
The true soft mode lies near 2/3 of $\Gamma-M$. At $M$, all modes are linearly polarized, but as one goes toward $\Gamma$, one set of atoms moves in the orthogonal direction to form an elliptical orbit. Notably, in 1H-NbSe2 an earlier work found two static CDW ground states, corresponding to the breathing-in and breathing-out patterns, with an energy difference of $\sim$8 meV.[NbSe2]
In the two longitudinal modes 3 and 9, the clustered phonon displacements are similar, differing only in the phase of the Se displacements with respect to Nb: when Nb moves toward the line of Se atoms, the Se atoms move toward (away from) the Nb plane in mode 9 (3), Fig. 8, so the change in eigenfunctions upon anticrossing amounts merely to flipping the Se phase. In the intermediate mode 6 the Se is rotated from near-Z to near-Y as it moves toward Z.
Figure 8: (color online) 2($\Gamma$-M)/3-point phonon displacements. (a) to (c) show modes 8, 6 and 3 respectively in the Y-Z plane, where, ${\bf Z\parallel a_{3}},$ ${\bf X\parallel a_{1}\parallel\Gamma-M}$, and ${\bf Y\perp X,Z}.$ Blue (orange) dots refers to Nb (Se) atoms.(d)Phonon dispersions. Red dots refer to the phonons considered.
We note that time molecules such as Na3 are chiral. However, a solid which is composed of many orbitals, can only be chiral if there is an imbalance between the clockwise and anticlockwise rotating molecules. Therefore, bulk NbSe2, which is centrosymmetric, cannot be chiral, but chirality is allowed in monolayer NbSe2.
VIII.3 Beyond the Born-Oppenheimer (BO) approximation
The EPW code, see Appendix B, includes a perturbative nonadiabatic (i.e., beyond BO) correction, which we have found to be small in NbSe2, although nonadiabatic effects can be much stronger.[nonadi, BBO1, BBO2]
Since the BO approximation relies on the large difference in the time scales involved in the electronic and ionic motions, nonadiabatic corrections would be larger when electrons move slowly as in heavy Fermion materials, flat bands, and near van Hove singularities (VHSs). Optical phonons can have strong nonadiabatic effects, as electron hoppings can be strongly enhanced for certain phonon motions, as discussed in the preceding subsection.
VIII.4 Estimating temperature
The broadening parameter $\sigma$ is often introduced as a fictitious temperature to improve the convergence of the DFT self-consistency cycles towards the ground state[fict]. Using the Fermi function with $\sigma=k_{B}T_{\sigma}$, however, leads to a large volume of $k$-space that must be integrated over, so that a more confined broadening function, such as a Gaussian, is a better choice. To gaing a handle on relating $\sigma$ to an effective temperature, we can compare the results of $\sigma$-broadening with direct calculations of the $T$-dependent Lindhard susceptibility, Fig. 9. Focusing on the broadening of the susceptibility peak along $\Gamma-M-K$, it is seen to be very small up to $\sigma=0.016$ Ry ($T_{\sigma}=2526$ K) where the peak is less reduced than for $T=250K$. But, the $\sigma=0.02$ susceptibility ($T_{\sigma}=3157$K) is much more intense and broader than at other $\sigma$ values, with a peak-width that is comparable to the $T=1000$K susceptibility. These results suggest that $T_{\sigma}/T$ is $\leq 10$ and $\geq 3.5$. Since the CDW transition occurs at $\sigma=0.0057$ Ry, or $T_{\sigma}=900$K, we thus estimate that the corresponding $T_{CDW}$ lies between 225K and 90K, which is consistent with the experimental value for monolayer NbSe2, $T_{CDW}=145$K.
Figure 9: (color online) Left panel: Susceptibility using the minimum $\sigma$ value at different Fermi-Dirac temperatures. Right panel: Susceptibility for fixed Fermi-Dirac temperature $T=1$ K for various $\sigma$ values.
IX Conclusions
We show that by accounting for effects of the divergent factor of $\sim 1/\Omega_{q}$ in the self-energy, new insight can be obtained into the nature of the phonon spectra of materials.
In this connection, we have incorporated in the EPW code the computation of off-diagonal mode-mode couplings in the phonon self-energy, which has enabled us to uncover the role of anti-crossing phenomenon in generating Kohn anomalies. An in-depth study of the Giant Kohn anomaly and chiral phonons in 1H-NbSe2 transition-metal-dichalcogenide is presented as an illustrative application of our modified EPW code. We provide compelling evidence that the phonon that softens in NbSe2 is a longitudinal optical phonon which softens by anti-crossing several intervening phonon bands. We also show that the CDW vector in NbSe2 results from the convolution of the susceptibility and electron-phonon coupling, and that the softened phonons are circularly polarized. These chiral phonons may be relevant for generating time crystals, reducing decoherence in solid-state defect-spin qubits, among other potential applications. Our study will impact the modeling of charge density waves and other phenomena involving electron-phonon couplings by enabling accurate, material-specific computation of phonon self-energies in wide classes of materials.
Appendix A Deconvolving the self-energy
To better understand the origin of the two peaks at $q_{1}$ and $q_{2}$, found in $\Pi$, we plot the $k$-resolved contributions to $\Pi(\Omega_{q})$ in Fig.A.1. Panel (a) shows electronic dispersion over the full Brillouin zone and the Fermi surface to determine corresponding vectors for the peaks. In panel (c), we plot $|g_{k,{k+q}}|^{2}$, in panel (b) $\chi_{q,k}$, the integrand of Eq. 7, and in panel (b) we superpose the two plots. We have confirmed that the integral of this superposition over the Brillouin zone gives the calculated values of $\Pi$ at those $q$-values. We note that the $\chi$ maps measure nesting – the overlap of $\epsilon_{k}$ and $\epsilon_{k+q}$, while the $G$ maps tend to enhance features in approximately the same $k$-regions of the susceptibility maps, where the nesting is significantly better for $q_{2}$. In Fig. A.2 we compare the clean and dirty limit self energies.
Figure A.1: (color online) Deconvolution of self-energy, (a)Left: Energy dispersion over the full Brillouin zone, with energy increasing from dark blue to dark red, Right: Fermi surface, (b)-(d) are respectively $G(k)\cdot\chi(k)$, $G(k)=|g_{\bf k,\bf k+\bf q}|^{2}$ and $\chi(k)$. Left figures are for peak between $\Gamma M$ (peak $q_{1}$) and Right ones are for the peak between $MK$ (peak $q_{2}$)
Figure A.2: (color online) Phonon spectra and self energies at three effective temperatures, from left to right $\sigma$ = 0.0057, 0.009, 0.02 Ry. Row (a) phonon modes, (b) self-energy in clean limit, and (c) self-energy in dirty limit. All frames use same color codes to identify the different phonon modes.
Appendix B Density functional perturbation theory (DFPT) and self-energies
The dressed phonon frequencies $\Omega_{q}$ are calculated in QE directly from density functional perturbation theory (DFPT) [Baroni]. To better understand the ‘bare’ phonon frequencies we have found, we briefly recapitulate the derivation of the QE formalism[QuantEspRenorm].
The physical basis of these calculations is the Born-Oppenheimer approximation, where the nuclear degrees of freedom can be separated from electronic degrees of freedom. Within the harmonic approximation, the phonon spectra and eigenmodes can be calculated using the Hessian of the BO-energy surface scaled by the geometric mean of the masses of the nuclei
$\left(\frac{1}{\sqrt{M_{I}M_{J}}}\frac{\partial^{2}E(R)}{\partial R_{I}\partial R_{J}}-\omega^{2}\mathbf{I}_{3N\times 3N}\right)\mathbf{\hat{u}}=0,$ (13)
where the potential energy surface $E(R)$ is calculated using the BO Hamiltonian for electrons, and $\mathbf{I}_{3N\times 3N}$ is the identity matrix. The non-trivial solutions give the eigenfrequencies $\omega_{\mu}$ and eigenvectors $\mathbf{\hat{u}}_{\mu}$ for phonons. In Baroni et al. [Baroni], the Hellman-Feynman theorem and BO-Hamiltonian are applied, which leads to a useful form of of Hessian:
[…]
where the $\mathbf{F}_{I}$ are the interatomic forces and $V_{R}(r)$ is the electron-nucleus interaction.
Equation 14 is the chief result of this Appendix. While not immediately obvious, the first term
[…]
contains a static coupling between electrons and phonons, since using linear response theory (see, e.g., Baroni Eq. (69) [Baroni]) the first order variation of the electron density in terms of Kohn-Sham eigenenergies and eigenfunctions indicates the appearance of a non-retarded (static) response function
[…]
When Fourier transformed and transformed into the basis of phonon modes, this corresponds to the static limit as Equation 9. Thus, the first term of Eq. 14 corresponds to our $\Pi^{\prime}$, so the remaining terms ${\bf\Pi}^{B}$ constitute $\omega_{q}^{2}$. We note that one term is the potential of the static ions, while the other is the interaction of the ions with the average electron density $\left\langle n_{R}\right\rangle$, which has a temperature dependence via the Fermi function. Moreover, since the standard DFpT calculation in QE includes the static part of the charge density variation, this term includes the main difference between $\Omega_{q}^{2}$ and the “bare” phonons $\omega_{q}^{2}$, using the definition of Ref. [Giustino1].
To further relate our results to current literature, we note the following: $\Pi^{\prime}_{\lambda,\lambda^{\prime}}(\boldsymbol{q})$ is the adiabatic self-energy as defined by Giustino [Giustino1]. If this is subtracted from the frequency dependent self-energy of Eq. 9 one obtains a non-adiabatic self-energy term with a retarded response of the electronic degrees of freedom to displacements of nuclei. In EPW the DFpT phonons are the bare phonons which are dressed by the non-adiabatic self energy to produce the dressed or dynamical phonons. More specifically, the correction is assumed to be small enough that only the diagonal self energy is important, so the non-adiabatic self energy is defined as $\Pi^{NA^{\prime}}(q,\Omega_{q})=\Pi^{\prime}_{\lambda,\lambda}({\boldsymbol{q}},\omega=\Omega_{\lambda})-\Pi^{\prime}_{\lambda,\lambda}(\boldsymbol{q})$ (Eq. 145 in Ref. Giustino1). In fact, adiabatic and non-adiabatic self-energy terms can be equated to static and dynamical terms of self-energy, as explicitly discussed in Ref. [ultrasoft], where they make a distinction between static and dynamical phonons ((bare and dressed phonons in the terminology of Ref. [Giustino1]).
However, we find that this correction is small, and the main softening is contained in the DFPT phonons, necessitating our rewriting of the dynamical matrix to find the bare BO phonons – i.e., without the e-ph coupling self energy.
Finally, we note that in Eq. 14 the unknown $\omega$ arises only in the $\omega^{2}$ term, inconsistent with Eq. 17, but consistent with Eq. 10. This supports our idea that in the term $\omega\Pi$, the prefactor $\omega$ exactly cancels the normalization-derived $1/\omega$ in $\Pi$, leaving $\Pi^{\prime}$ the essential quantity to calculate.
Appendix C Inverse dynamical matrix
In trying to find the bare phonon frequencies in terms of the dressed phonons, there are two possible choices. The simplest procedure is to modify Eq. 1 by replacing $\omega_{\boldsymbol{q}}$ by $\Omega_{\boldsymbol{q}}$ in the self-energy $\Pi$,
$\det\Bigl{\{}({\boldsymbol{\Omega}}_{\boldsymbol{q}}^{2}-{\omega^{2}})\operatorname{\boldsymbol{I}}-2\,{\omega}\,{\boldsymbol{\Pi}}(\boldsymbol{q},{\boldsymbol{\Omega}}_{\boldsymbol{q}})\Bigr{\}}=0.$ (17)
Since the unknown $\omega$ occurs explicitly in the dynamical matrix, the bare frequencies are not the solutions of a determinant equation, but must be found self-consistently. However, if the off-diagonal elements of $\Pi$ are neglected, Eq. 17 becomes a quadratic equation. The inverse solution to Eq. 17 is
$\omega_{q}=-\Pi(\Omega_{q})+\sqrt{\Omega_{q}^{2}+\Pi^{2}(\Omega_{q})}\rightarrow-2\Pi(\Omega_{q}),$ (18)
where the limiting behavior holds near the soft mode, $\Omega_{q}=0$ (recall that $\Pi$ is negative). Thus, Eq. 17 leads to an unphysical divergence of $\omega_{q}$ every time there is a soft mode, Fig. C.1. We believe that this is a shortcoming of the Hedin approach, in that for an equation like Eq. 1 it is not clear which $\omega_{q}$-terms remain, and which are replaced by $\Omega_{q}$. Here, we propose an alternative solution. Noting that in the term $\Pi^{\prime}(\omega_{q})=\omega_{q}\Pi(\omega_{q})$, the prefactor $\omega_{q}$ exactly cancels the normalization factor $1/\omega_{q}$ in $\Pi$, we propose that this should be an exact result. Equivalently, in the Hedin-like expression $\Pi^{\prime}(\omega_{q})$ should be replaced by $\Pi^{\prime}(\Omega_{q})$. When this is done, Eq. 17 becomes
$\det\Bigl{\{}(\Omega^{2}-{\boldsymbol{\omega}_{\boldsymbol{q}}^{2}})\operatorname{\boldsymbol{I}}-2\Omega{\boldsymbol{\Pi}}(\boldsymbol{q},\Omega)\Bigr{\}}=0,$ (19)
which can be reorganized to Eq. 10. Note that Eq. 10 can be thought of as the equation for the zeroes of an anti-Green’s function which derives the bare frequencies from the dressed. Neglecting off-diagonal self energies, the solution becomes
$\omega_{q}^{2}=\Omega_{q}^{2}-2\Omega_{q}\Pi(\Omega_{q})\rightarrow-2\Pi^{\prime}(\Omega_{q}).$ (20)
In Fig. 4 we show the bare phonon frequencies calculated from Eq. 20 as dashed lines. These results should be contrasted with Fig. C.1.
Figure C.1: (color online) Single mode calculations for the three longitudinal modes, modes 3, 6, and 8. Solution of Eq. 18 for $\sigma$ =0.057 Ry, comparing the dressed frequencies $\Omega_{q}$ (dashed curves) with the bare frequencies $\omega_{q}$ (solid curves).
Appendix D DFT Parameters and Potentials
Density functional theory (DFT) calculations were performed using the Quantum ESPRESSO package. The structural optimization was carried out for mono-layer 1H-NbSe2 by including ${\AA}$ and setting the energy and force convergence thresholds to $10^{-6}$ and $10^{-5}$ Ry, respectively. A plane-wave basis set was employed, with a wavefunction cutoff of 100 Ry and a charge density cutoff of 800 Ry. For the electronic structure, a convergence threshold of $10^{-12}$ Ry was used, incorporating local Thomas-Fermi screening. Fermi-Dirac smearing was applied to model temperature effects, and the smearing value was varied accordingly. Ultrasoft pseudopotentials were used for Nb, while norm-conserving pseudopotentials were selected for Se, following the methodology outlined in Ref. anharm. Self-consistent field (SCF) calculations were performed using a $36\times 36\times 1$ Monkhorst-Pack $K$-grid. Phonon spectra were computed with a precision of $10^{-16}$ Ry on a $12\times 12\times 1$ mesh. The electron-phonon coupling (EPC) matrix was evaluated using the EPW code along a $Q$-path sampled with 500 points, utilizing a dense $500\times 500\times 1$ $k$-point mesh.
Appendix E Kekule-like Graphene
Figure E.1: (color online) K-point phonon displacements in graphene. Top and bottom rows show the same three phases of a single cycle of the highest K-point phonon in graphene.
Using Phonon Visualizer, we find similar Kekule-like phonon displacement patterns in benzene, Fig. E.1. The top row focuses on a breathing-in mode for a single hexagon [yellow squares], as it cycles among three possible centers for the hexagon. In the bottom row, we focus instead on a single hexagon [purple squares] as it evolves through the same cycle. We observe the same Kekule-like pattern as seen in NbSe2.
Acknowledgements
The work at Northeastern University was supported by the National Science Foundation through the Expand-QISE award NSF-OMA-2329067 and benefited from the resources of Northeastern University’s Advanced Scientific Computation Center, the Discovery Cluster, the Massachusetts Technology Collaborative award MTC-22032, and the Quantum Materials and Sensing Institute. S.C. and S.E. acknowledge the support provided by the National Science Foundation through the ExpandQISE award No. 2329087. S.C. and S.M. are supported by the U.S. Department of Energy (DOE), Office of Science, and Basic Energy Sciences under Award No. DE-SC0022216. The calculations performed by T.H. and C.S. in this paper were performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources). The research at Howard University used the resources of Accelerate ACCESS PHYS220127 and PHYS2100073. We thank Adrian Feiguin for stimulating discussions. T.H. is thankful to Cagliyan Kurdak and the Department of Physics of the University of Michigan for partial support.